# Supermassive black holes from inflation constrained by dark matter substructure

### v1

@@ I Introduction
The James Webb Space Telescope (JWST) is revealing the high-redshift universe in novel ways, offering much more insight into the seeds of SMBHs by studying their high redshift population. For instance, there is evidence for SMBHs at $118<z<$ 11 with candidates in the range $M_{\textrm{BH}}={\cal O}(10^{6}$$10^{8})M_{\odot}$ Larson et al. (2023); Maiolino et al. (2023); Bogdan et al. (2024); Natarajan et al. (2024); Harikane et al. (2023). These observations mount further pressure on models with low mass progenitors without super-Eddington accretion or rapid merger rates in the early universe. Feedback from super-Eddington growth can hinder SMBH growth — jets and outflows from accretion push material away, so the duration of super-Eddington growth phases should be finite Massonneau et al. (2023) but a realistic value for the duration is still unknown. Merger scenarios for the SMBH origin require highly clustered early populations, which can affect SMBH growth both positively and negatively, as they can act to eject SMBHs from galaxy centers Banik et al. (2019).
@@
The homogeneity and isotropy of cosmic microwave background (CMB) observations impose strict constraints on the amplitude $A_{s}$ and spectral index $n_{s}$ of scalar perturbations. For example, at the pivot scale $k_{*}=0.05\leavevmode\nobreak\ {\rm Mpc}^{-1}$, The amplitude and the spectral index is constrained as $A_{s}=(2.099\pm 0.029)\times 10^{-9}$ and $n_{s}=0.9649\pm 0.0042$ (Planck TT,TE,EE+lowE data Aghanim et al. (2020)). At smaller scales, constraints on the curvature perturbations are much less stringent. Currently, the amplitude for comoving wavenumbers $k>\mathcal{O}({1})\leavevmode\nobreak\ {\rm Mpc}^{-1}$ is constrained through considerations of $\mu$- and $y$-type distortions in CMB observations Chluba et al. (2012, 2015), the overproduction of PBHs Josan et al. (2009); Carr et al. (2010); Byrnes et al. (2019); Dalianis (2019); Sato-Polito et al. (2019); Gow et al. (2021), density profiles of ultracompact minihalos Delos et al. (2018); Nakama et al. (2018), free-free emission in the Planck foreground analysis Abe et al. (2022), the galaxy luminosity function Yoshiura et al. (2020) and gravitational lensing Gilman et al. (2022). The enhancement of the primordial power spectrum of curvature perturbations $P_{\mathcal{\zeta}}$ at large $k\simeq{\cal O}(10^{2}$$10^{4})\,\textrm{Mpc}^{-1}$ is constrained to be lower than around ${\cal O}(10^{-4})$  [^1] The curvature perturbation in comoving coordinates $\zeta$ is used interchangeably with the comoving curvature perturbation $\mathcal{R}$ in the literature because they describe the same physical quantity during inflation and on superhorizon scales..
However, the amplitude of the primordial curvature perturbation needed to explain the abundance of PBHs as SMBH seeds can be decreased if the density distribution follows non-Gaussian statistics  Ünal et al. (2021). If this is the case, the current constraints listed above can be avoided.
@@ II Primordial curvature perturbations and PBH abundance
Power spectrum. We investigate models with a prominent feature at small scales with wavenumber $k\gtrsim\mathcal{O}({1})\leavevmode\nobreak\ {\rm Mpc}^{-1}$ in the primordial power spectrum. To evaluate the impact of the curvature perturbation in this range, we add an additional feature, namely an extra bump, on top of the nearly scale-invariant curvature perturbation spectrum that matches the features seen in the CMB
@@
$P_{\zeta}^{\rm bump}(k;k_{p})=\left\{\begin{array}[]{ll}(A-P_{\zeta}^{(0)}(k_{p}))\left(\frac{k}{k_{p}}\right)^{n_{b}}&k\leq k_{p}\\ 0&k>k_{p}\end{array}\right.\,.$ (4)
Here, we introduce three parameters denoted as $A$, $k_{p}$, and $n_{b}$, which describe the amplitude, the corresponding wavenumber, and the growth index of the bump, respectively. Considering single-field inflation models, Ref. Byrnes et al. (2019) suggests a steepest spectral index of $n_{b}=4$. Conversely, Ref. Ozsoy and Tasinato (2020) argues that the spectral index could reach as high as $8$ after experiencing a dip in amplitude and subsequently peaking with an index less than 4. We show some illustrative examples of the primordial power spectrum with different amplitudes and growth indices as a function of the ratio $k/k_{p}$ in Fig. 1.
@@
While the mass of the PBH at formation is described using the comoving wavenumber of the fluctuation Nakama et al. (2017a), its evolution due to accretion and mergers can be parameterized by introducing two parameters ${\cal A}$ for accretion and ${\cal M}$ for mergers Garcia-Bellido et al. (2017)
@@
$\beta_{\rm G}$ $={\rm erfc}\left(\frac{\zeta_{c}}{\sqrt{2P_{\zeta}/{\cal K}}},\right),$
$\ \beta_{\chi^{2}}$ $={\rm erfc}\left(\sqrt{\frac{1}{2}+\frac{\zeta_{c}}{\sqrt{2P_{\zeta}/{\cal K}}}},\right),$
$\beta_{{\rm G}^{3}}$ $={\rm erfc}\left[\left(\frac{\zeta_{c}}{\sqrt{8P_{\zeta}/15{\cal K}}}\right)^{1/3}\right].$ (10)
We note that $\zeta_{c}$ is the critical value of the curvature perturbation which represents the deviation from flatness in a region. Nonlinearities in the Press-Schechter or peak theory formalism are accounted for by introducing the factor ${\cal K}$ in the standard threshold integral. In the following calculation, ${\cal K}$ in Eq. (10) is fixed to 2. We note that mode coupling makes it hard to keep a strict $\chi^{2}$ distribution at small scales Young and Byrnes (2015); but these uncertainties are negligible for our cases considering narrow spectra. For the following, we will consider when $\zeta_{c}=\delta_{c}=0.45$ Carr (1975); Musco and Miller (2013), as this is the condition required of the density contrast to form a black hole.
@@
Figure 2: Dependence of the initial fraction of causal horizons $\beta$ as a function of the primordial power spectrum amplitude. The Gaussian case is shown as a cyan (dot-dashed), the $\chi^{2}$ case as a dark cyan (dashed), and the ${\rm G}^{3}$ case as a blue (solid) line. The plots are corresponding to the case assuming $\mathcal{K}=2$ and $\zeta_{c}=\delta_{c}=0.45$
We relate $\beta$ to the current energy density in PBHs of seed mass $M_{\textrm{seed}}$ as Garcia-Bellido et al. (2017); Ünal et al. (2021)
[…]
@@
This quantity is to be compared to the energy density of SMBHs, $\Omega_{\rm SMBH}$. Ref. Ünal et al. (2021) evaluates $\Omega_{\rm SMBH}$ from the number of the galaxies that can host massive black holes Conselice et al. (2016) and the mass range of SMBHs as $10^{6}$$10^{8}M_{\odot}$, obtaining $10^{-7}\lesssim\Omega_{\rm SMBH}\lesssim 10^{-5}$ by dividing the sum of the mass in those SMBH by the total mass of the Universe. Following the estimate in the previous work, we set $\Omega_{{\rm PBH},0}=10^{-10}$ as a conservative benchmark to test the PBH origin scenario for SMBHs.
@@ III Dark matter subhalos
In the EPS theory, instead of halo mass and redshift $(m,z)$, one adopts $(S,\delta_{c}^{\textrm{ha}})$. $\delta_{c}^{\textrm{ha}}$ is the halo collapse threshold at redshift $z$ in the linearly extrapolated spherical collapse model, $\delta_{c}^{\textrm{ha}}\equiv 1.686/D(z)$, where $D(z)$ is the linear growth factor. $S\equiv\sigma^{2}(m)$ is the variance of the density fluctuation which we evaluated as that at $z=0$, where we apply the sharp-$k$ filter  [^2] The filter is suitable for power spectra with a steep cutoff Schneider et al. (2013).. The mass fraction contained in smaller halos which collapse at redshift $z_{2}$, where $\delta^{\textrm{ha}}(z_{2})$ is written $\delta^{\textrm{ha}}_{2}$, with corresponding mass scale $S_{2}$, characterized by $(S_{2},\delta_{2}^{\textrm{ha}})$ in their host halo $(S_{1},\delta_{1}^{\textrm{ha}})$ of $z_{1}<z_{2}$ is given by
@@
×$\times\exp\left[-\frac{(\delta_{2}^{\textrm{ha}}-\delta_{1}^{\textrm{ha}})^{2}}{2(S_{2}-S_{1})}\right]dS_{2}.$
@@ IV Results and discussion
In Fig. 3, the prediction of the subhalo mass function with several values of $A$ are plotted in different lines. The left (right) panel corresponds to the case of $n_{b}=4$ ($n_{b}=8$). We fix the parameter $k_{p}=32\,h\,\mathrm{Mpc^{-1}}$ where $h=0.6736$ Aghanim et al. (2020). The figure clearly shows that the subhalo mass function within the vicinity of the Milky Way is affected by the injection of power on this characteristic scale. For the same wavenumber, the maximum in the subhalo mass function shifts to higher masses as the amplitude of the power enhancement increases. The number of smaller subhalos, on the other hand, is suppressed because halos corresponding to wavenumbers of the power spectrum bump collapse so early that smaller substructures cannot form as they will be embedded inside these larger halos.
@@
Figure 3: The subhalo mass function $dN_{\rm sh}/d\ln m$ within 300 kpc volume of the Milky-Way halo. The data points are a combination of the stellar-stream and lensing measurements Banik et al. (2021a, b). The curves are predicted subhalo mass functions for $k_{p}=32\,h$ Mpc-1 and smaller to larger values of $A$ (from top to bottom). The left and right panels are for $n_{b}=4$ and $n_{b}=8$, respectively. The dotted line in each panel corresponds to the prediction without a bump feature.
@@
The Dark Energy Survey and PanSTARRS1 survey identified 94 satellites with the kinematic data within the virial volume of the Milky Way after imposing the completeness correction Drlica-Wagner et al. (2020); Dekker et al. (2021). The satellite galaxy with the smallest velocity dispersion, $\sigma_{V}$, is Leo V with $\sigma_{V}=2.3$ km s-1 Drlica-Wagner et al. (2020). We take this value as the minimum for subhalos to host galaxies, and scan the parameter region where the number of subhalos satisfying the $V_{\rm max}^{\rm Leo\leavevmode\nobreak\ I}=\sqrt{3}\sigma_{V}=4$ km s-1 is larger than 94.
@@
In the example shown in Fig. 3, we see that the model with e.g. $A=10^{-2}$ predicts too small numbers of subhalos in the mass range where lensing and stellar stream observations are sensitive, hence it can be excluded. We will also discuss future projections in the following sections.
@@
Figure 4: Constraints from current observations on the peak amplitude of the primordial power spectrum $P_{\zeta}$ as a function of the peak wavenumber $k_{p}$ and PBH seed mass. The left panel assumes $n_{b}=4$ and the right panel does $n_{b}=8$. The light and dark solid purple lines represent constraints on the power spectrum from the satellite counts, and combined stream and lensing analysis at the 95% confidence level, respectively. The solid gray line corresponds to COBE/FIRAS $\mu$-distortion bounds. The black region is excluded by Lyman-$\alpha$ data Bird et al. (2011). The solid cyan, dark cyan, and blue regions correspond to the magnitude of Gaussian, $\chi^{2}$, and ${\rm G}^{3}$ distributions which would give rise to an SMBH density of $\Omega_{\textrm{SMBH}}=10^{-10}$, respectively. The upper and lower boundary of these three bands correspond to the accretion factors $\mathcal{A}=1$ and $\mathcal{A}=10^{5}$.
Sensitivity to the possible SMBH seeds from current observations are summarized in Fig. 4. The left (right) panel shows the case which assumes $n_{b}=4$ ($n_{b}=8$). In each panel, we plot the excluded region from satellite number counts with light purple. The constraints obtained from lensing and stellar stream observations, which are shown in dark purple, surpass those from satellite number counts. By combining the lensing and stellar stream observations, the sensitivity improves by a factor of $\simeq 10$. If the parameters $(A,k_{p})$ are above these lines, our model yields fewer than 94 subhalos with $V_{\rm max}>4$ km s-1, hence an upper limit at the 95% confidence level is obtained. The sensitivity with subhalo number counts reaches as small as $P_{\zeta}\sim 3\times 10^{-7}$ at $k_{p}\sim{\cal O}(10)\,{\rm Mpc}^{-1}$.
Figure 5: Future prospects on the sensitivity to the peak amplitude of the primordial power spectrum $P_{\zeta}$. Projected constraints from galaxy stream observations are plotted in solid magenta lines corresponding to less than a 100% increase in subhalos ($\frac{N_{\rm sh}(A,k_{p},n_{b})}{N_{\rm sh}(A=0)}<2$) in four subhalo mass ranges, $10^{5}$$10^{6}$, $10^{6}$$10^{7}$, $10^{7}$$10^{8}$, and $10^{8}$$10^{9}M_{\odot}$, denoted using the lightest to darkest magenta colours respectively. Here we assume no uncertainty in the subhalo mass function for the scale-invariant power spectrum. Prospects with future $\mu$-distortion measurements with PIXIE Ünal et al. (2021) are shown with gray-dashed lines. See Fig. 4 for the Lyman-$\alpha$ bounds and expected amplitude for curvature perturbations to generate SMBH seeds.
@@
Here we denote ${\hat{k}}=k\,/\,{\rm Mpc}$ and ${\hat{k}}_{0}=1$. Current CMB $\mu$-distortion bounds are sensitive to the amplitude down to $P_{\zeta}\sim 3\times 10^{-3}$. The probe is complementary to those DM substructures, which can be more sensitive in the mass range of $M_{\rm PBH}^{\rm seed}\gtrsim{\cal O}(10^{9})M_{\odot}$.
In each panel, the amplitude of the primordial fluctuation which produces the current SMBH energy density of $\Omega_{\textrm{SMBH}}=10^{-10}$ assuming three different statistics, $\beta_{\rm G},\beta_{\chi^{2}}$ and $\beta_{{\rm G}^{3}}$nal et al. (2021) are shown with light cyan, dark cyan, and blue, respectively. Depending on the accretion efficiency ${\cal A}$, the amplitude satisfying the SMBH abundance can be in between the upper (${\cal A}=1$) or lower (${\cal A}=10^{5}$) boundaries of the bands.
The sensitivity with DM structure formation is sufficient to probe non-Gaussian statistics of $\chi^{2}$ and $G^{3}$ distributions at scales $k_{p}\lesssim{\cal O}(10^{2}){\rm Mpc}^{-1}$. For $n_{b}=8$, the accessible region gets slightly narrower, however, up to $k_{p}\sim 80\,{\rm Mpc}^{-1}$ can still be probed. These regions correspond to PBH seed masses $\gtrsim 10^{9}M_{\odot}$. From the measurements of the CMB $\mu$-distortion by COBE/FIRAS, the Gaussian statistics case over scales of $10^{1}\lesssim k_{p}\cdot\textrm{Mpc}\lesssim 3\times 10^{4}$ for both $n_{b}$ are already excluded. For a narrower range of $10^{2}\lesssim k_{p}\cdot\textrm{Mpc}\lesssim 10^{4}$, COBE/FIRAS observations exclude the Gaussian case if $n_{b}=4$ while it does not for $n_{b}=8$. In all cases, the most non-Gaussian (${\rm G}^{3}$) case is not yet explored by $\mu$-distortion measurements.

### v3

@@ I Introduction
The James Webb Space Telescope (JWST) is revealing the high-redshift universe in novel ways, offering much more insight into the seeds of SMBHs by studying their high redshift population. For instance, there is evidence for SMBHs at $8<z<11$ with candidates in the range $M_{\textrm{BH}}={\cal O}(10^{6}$$10^{8})M_{\odot}$ Larson et al. (2023); Maiolino et al. (2023); Bogdan et al. (2024); Natarajan et al. (2024); Harikane et al. (2023). These observations mount further pressure on models with low mass progenitors without super-Eddington accretion or rapid merger rates in the early universe. Feedback from super-Eddington growth can hinder SMBH growth — jets and outflows from accretion push material away, so the duration of super-Eddington growth phases should be finite Massonneau et al. (2023) but a realistic value for the duration is still unknown. Merger scenarios for the SMBH origin require highly clustered early populations, which can affect SMBH growth both positively and negatively, as they can act to eject SMBHs from galaxy centers Banik et al. (2019).
@@
The homogeneity and isotropy of cosmic microwave background (CMB) observations impose strict constraints on the amplitude $A_{s}$ and spectral index $n_{s}$ of scalar perturbations. For example, at the pivot scale $k_{*}=0.05\penalty 10000\ {\rm Mpc}^{-1}$, The amplitude and the spectral index is constrained as $A_{s}=(2.099\pm 0.029)\times 10^{-9}$ and $n_{s}=0.9649\pm 0.0042$ (Planck TT,TE,EE+lowE data Aghanim et al. (2020)). At smaller scales, constraints on the curvature perturbations are much less stringent. Currently, the amplitude for comoving wavenumbers $k>\mathcal{O}({1})\penalty 10000\ {\rm Mpc}^{-1}$ is constrained through considerations of $\mu$- and $y$-type distortions in CMB observations Chluba et al. (2012, 2015), the overproduction of PBHs Josan et al. (2009); Carr et al. (2010); Byrnes et al. (2019); Dalianis (2019); Sato-Polito et al. (2019); Gow et al. (2021), density profiles of ultracompact minihalos Delos et al. (2018); Nakama et al. (2018), free-free emission in the Planck foreground analysis Abe et al. (2022), the galaxy luminosity function Yoshiura et al. (2020) and gravitational lensing Gilman et al. (2022). The enhancement of the primordial power spectrum of curvature perturbations $P_{\mathcal{\zeta}}$ at large $k\simeq{\cal O}(10^{2}$$10^{4})\,\textrm{Mpc}^{-1}$ is constrained to be lower than around ${\cal O}(10^{-4})$  [^1] The curvature perturbation in comoving coordinates $\zeta$ is used interchangeably with the comoving curvature perturbation $\mathcal{R}$ in the literature because they describe the same physical quantity during inflation and on superhorizon scales..
However, the amplitude of the primordial curvature perturbation needed to explain the abundance of PBHs as SMBH seeds can be decreased if the density distribution follows non-Gaussian statistics Ünal et al. (2021); Iovino et al. (2024). If this is the case, the current constraints listed above can be avoided.
@@ II Primordial curvature perturbations and PBH abundance
Power spectrum. We investigate models with a prominent feature at small scales with wavenumber $k\gtrsim\mathcal{O}({1})\penalty 10000\ {\rm Mpc}^{-1}$ in the primordial power spectrum. To evaluate the impact of the curvature perturbation in this range, we add an additional feature, namely an extra bump, on top of the nearly scale-invariant curvature perturbation spectrum that matches the features seen in the CMB
@@
$P_{\zeta}^{\rm bump}(k;k_{p})=\left\{\begin{array}[]{ll}n_{b}(A-P_{\zeta}^{(0)}(k_{p}))\left(\frac{k}{k_{p}}\right)^{n_{b}}&k\leq k_{p}\\ 0&k>k_{p}\end{array}\right.\,$. (4)
Here, we introduce three parameters denoted as $A$, $k_{p}$, and $n_{b}$, which describe the amplitude, the corresponding wavenumber, and the growth index of the bump, respectively. Note the coefficent $n_{b}$ in Eq. (4) is necessary to ensure that the variance of the primordial power spectrum is normalized to $A$ as per convention. Considering single-field inflation models, Ref. Byrnes et al. (2019) suggests a steepest spectral index of $n_{b}=4$. Conversely, Ref. Ozsoy and Tasinato (2020) argues that the spectral index could reach as high as $8$ after experiencing a dip in amplitude and subsequently peaking with an index less than 4. We show some illustrative examples of the primordial power spectrum with different amplitudes and growth indices as a function of the ratio $k/k_{p}$ in Fig. 1.
@@
[…]
@@
$\beta_{\rm G}$ $={\rm erfc}\left(\frac{\zeta_{c}}{\sqrt{2}\sigma/\sqrt{\mathcal{K}}}\right),$
$\ \beta_{\chi^{2}}$ $={\rm erfc}\left(\sqrt{\frac{1}{2}+\frac{\zeta_{c}}{\sqrt{2/{\cal K}}\sigma}}\right),$
$\beta_{{\rm G}^{3}}$ $={\rm erfc}\left[\left(\frac{\zeta_{c}}{\sqrt{8/{15\cal K})}\sigma}\right)^{1/3}\right].$ (10)
We evaluate the variance $\sigma$ that enters Eq. (10) as a function of the smoothing scale $R=k^{-1}$
$\sigma^{2}(k)=\frac{16}{81}\int d\ln k^{\prime}\,\mathcal{P}_{\zeta}(k^{\prime})\left(\frac{k^{\prime}}{k}\right)^{4}\,|W(k^{\prime}/k)|^{2}\,T^{2}(k^{\prime}/k)\,.$ (11)
Here we adopt the top–hat window function $W(x)=3(\sin x-x\cos x)/x^{3}$ and the radiation–era transfer function $T(x)=3(x/\sqrt{3})^{-3}[\sin(x/\sqrt{3})-(x/\sqrt{3})\cos(x/\sqrt{3})]$, as in Ref. Kawasaki and Nakatsuka (2019). For a narrow peak in the primordial power spectrum, $\sigma^{2}(R)$ attains its maximum when $k_{p}R\simeq 2.5$33, as shown in Ref. Kawasaki and Nakatsuka (2019). [^2] It is pointed out in Ref. Ando et al. (2018) that the choice of the window function causes large uncertainties in the evaluation of the variance. For instance, we have checked that if we use the gaussian window function, then the resultant amplitude $A$ to give $\Omega_{\rm PBH,0}^{\rm tot}=10^{-10}$, defined in Eq. (13), becomes more than one order of magnitude larger.
When computing the variance, expressed as an integral over the external wavenumber $k$, we therefore evaluate the variance at the corresponding smoothing scale $R$ associated with that mode. This ensures that the physically relevant collapse scale is automatically incorporated by the window function and transfer function in Eq. (11).
We note that $\zeta_{c}$ is the critical value of the curvature perturbation which represents the deviation from flatness in a region. Nonlinearities in the Press-Schechter or peak theory formalism are accounted for by introducing the factor $\cal K$} in the standard threshold integral. The factor $\mathcal{K}$ in the Press-Schechter formalism is set to 2 to correct for the underestimation of the mass fraction of collapsed objects by accounting for both overdense and underdense regions. This adjustment ensures that the predicted mass fraction aligns with physical expectations, effectively capturing the overall impact of non-linearities in a statistical sense. We note that mode coupling makes it hard to keep a strict $\chi^{2}$ distribution at small scales Young and Byrnes (2015); but these uncertainties are negligible for our cases considering narrow spectra. For the following, we will consider when $\zeta_{c}=\delta_{c}=0.45$ Carr (1975); Musco and Miller (2013), as this is the condition required for the density contrast to form a black hole.
@@
Figure 2: Dependence of the initial fraction of causal horizons $\beta$ as a function of the primordial power spectrum amplitude. The Gaussian case is shown as a cyan (dot-dashed), the $\chi^{2}$ case as a dark cyan (dashed), and the ${\rm G}^{3}$ case as a blue (solid) line. The plots are corresponding to the case assuming $\mathcal{K}=2$, $\zeta_{c}=\delta_{c}=0.45$
We relate $\beta$ to the current energy density in PBHs of seed mass $M_{\textrm{seed}}$ (described conventionally by logarithmic intervals in seed mass) as Garcia-Bellido et al. (2017); Ünal et al. (2021)
$\Omega_{\rm PBH,0}$ $=\frac{\rho_{\rm PBH,0}(M_{\textrm{seed}},{\cal A},{\cal M})}{\rho_{\rm crit,0}}\newline$
@@
This quantity is to be compared to the energy density of SMBHs, $\Omega_{\rm SMBH}$. Ref. Ünal et al. (2021) evaluates $\Omega_{\rm SMBH}$ from the number of the galaxies that can host massive black holes Conselice et al. (2016) and the mass range of SMBHs as $10^{6}$$10^{8}M_{\odot}$, obtaining $10^{-7}\lesssim\Omega_{\rm SMBH}\lesssim 10^{-5}$ by dividing the sum of the mass in those SMBH by the total mass of the Universe. Following the estimate in the previous work, we set $\Omega^{\rm tot}_{\rm PBH,0}=10^{-10}$ as a conservative benchmark to test the PBH origin scenario for SMBHs.
@@ III Dark matter subhalos
In the EPS theory, instead of halo mass and redshift $(m,z)$, one adopts $(S,\delta_{c}^{\textrm{ha}})$. $\delta_{c}^{\textrm{ha}}$ is the halo collapse threshold at redshift $z$ in the linearly extrapolated spherical collapse model, $\delta_{c}^{\textrm{ha}}\equiv 1.686/D(z)$, where $D(z)$ is the linear growth factor. $S\equiv\sigma^{2}(m)$ is the variance of the density fluctuation which we evaluated as that at $z=0$, where we apply the sharp-$k$ filter. The filter is suitable for power spectra with a steep cutoff Schneider et al. (2013). The mass fraction contained in smaller halos which collapse at redshift $z_{2}$, where $\delta^{\textrm{ha}}(z_{2})$ is written $\delta^{\textrm{ha}}_{2}$, with corresponding mass scale $S_{2}$, characterized by $(S_{2},\delta_{2}^{\textrm{ha}})$ in their host halo $(S_{1},\delta_{1}^{\textrm{ha}})$ of $z_{1}<z_{2}$ is given by
@@
$\times\exp\left[-\frac{(\delta_{2}^{\textrm{ha}}-\delta_{1}^{\textrm{ha}})^{2}}{2(S_{2}-S_{1})}\right]dS_{2}.$
@@ IV Results and discussion
In Fig. 3, the prediction of the subhalo mass function with several values of $A$ are plotted in different lines. The left (right) panel corresponds to the case of $n_{b}=4$ ($n_{b}=8$). We fix the parameter $k_{p}=32\,h\,\mathrm{Mpc^{-1}}$ where $h=0.6736$ Aghanim et al. (2020). In the calculation for $n_{b}=8$ case, we use the host halo mass evolution with $n_{b}=4$ to save computational time, since the impact is expected to be small. The figure clearly shows that the subhalo mass function within the vicinity of the Milky Way is affected by the injection of power on this characteristic scale. For the same wavenumber, the maximum in the subhalo mass function shifts to higher masses as the amplitude of the power enhancement increases. The number of smaller subhalos, on the other hand, is suppressed because halos corresponding to wavenumbers of the power spectrum bump collapse so early that smaller substructures cannot form as they will be embedded inside these larger halos.
This behavior is qualitatively consistent with numerical simulations, such as those presented in Ref. Delos et al. (2018). In particular, Fig. 4 of Ref. Delos et al. (2018) demonstrates that a sharp enhancement in the power spectrum results in a significant abundance of halos at that corresponding scale, while suppressing the formation of smaller structures. These trends are in good agreement with our findings based on the EPS formalism. While a detailed quantitative comparison between our semi-analytical predictions and numerical simulations remains a task for future work, preliminary comparisons with N-body simulations indicate promising agreement, supporting the validity of our model.
We also note that the choice of a filter function (e.g., sharp-$k$ or top-hat) can introduce some ambiguity when comparing different power spectrum models. However, Ref. Ando et al. (2022) demonstrated that this ambiguity has minimal impact in cases with a sharply peaked power spectrum, supporting the reliability of our predictions.
@@
Figure 3: The subhalo mass function $dN_{\rm sh}/d\log_{10}m$ within 300 kpc volume of the Milky-Way halo. The data points are a combination of the stellar-stream and lensing measurements Banik et al. (2021a, b). The curves are predicted subhalo mass functions for $k_{p}=32\,h$ Mpc-1 and smaller to larger values of AA (from top to bottom). The left and right panels are for $n_{b}=4$ and $n_{b}=8$, respectively. The dotted line in each panel corresponds to the prediction without a bump feature. Note that the peak amplitude in the legend is simply $A_{\textrm{peak}}=A\,n_{b}$.
@@
The Dark Energy Survey and PanSTARRS1 survey identified 94 satellites with the kinematic data within the virial volume of the Milky Way after imposing the completeness correction Drlica-Wagner et al. (2020); Dekker et al. (2021). The satellite galaxy with the smallest velocity dispersion, $\sigma_{V}$, is Leo V with $\sigma_{V}=2.3$ km s-1 Drlica-Wagner et al. (2020). We take this value as the minimum for subhalos to host galaxies, and scan the parameter region where the number of subhalos satisfying the $V_{\rm max}^{\rm Leo\penalty 10000\ I}=\sqrt{3}\sigma_{V}=4$ km s-1 is larger than 94.
@@
In the example shown in Fig. 3, we see that the model with e.g. $An_{b}=10^{-2}$ predicts too small numbers of subhalos in the mass range where lensing and stellar stream observations are sensitive, hence it can be excluded. We will also discuss future projections in the following sections.
@@
Figure 4: Constraints from current observations on the amplitude of the primordial power spectrum AA as a function of the peak wavenumber $k_{p}$ and PBH seed mass. The left panel assumes $n_{b}=4$ and the right panel does $n_{b}=8$. The light and dark solid purple lines represent constraints on the power spectrum from the satellite counts, and combined stream and lensing analysis at the 95% confidence level, respectively. The solid gray line corresponds to COBE/FIRAS $\mu$-distortion bounds. The black region is excluded by Lyman-$\alpha$ data Bird et al. (2011). The solid cyan, dark cyan, and blue regions correspond to the magnitude of Gaussian, $\chi^{2}$, and ${\rm G}^{3}$ distributions which would give rise to an SMBH density of $\Omega_{\textrm{SMBH}}=10^{-10}$, respectively. The upper and lower boundary of these three bands correspond to the accretion factors $\mathcal{A}=1$ and $\mathcal{A}=10^{5}$.
Sensitivity to the possible SMBH seeds from current observations are summarized in Fig. 4. The left (right) panel shows the case which assumes $n_{b}=4$ ($n_{b}=8$). In each panel, we plot the excluded region from satellite number counts with light purple. The constraints obtained from lensing and stellar stream observations, which are shown in dark purple, surpass those from satellite number counts. By combining the lensing and stellar stream observations, the sensitivity improves by a factor of $\simeq 10$. If the parameters $(A,k_{p})$ are above these lines, our model yields fewer than 94 subhalos with $V_{\rm max}>4$ km s-1, hence an upper limit at the 95% confidence level is obtained. The sensitivity with subhalo number counts reaches as small as $A\sim 8\times 10^{-8}$ at $k_{p}\sim{\cal O}(10)\,{\rm Mpc}^{-1}$.
For the Lyman-$\alpha$ exclusion shown as a dark gray region, we evaluate the full primordial power spectrum at the characteristic Lyman-$\alpha$ scale $k_{\mathrm{Ly\alpha}}\simeq 2.7\penalty 10000\ {\rm Mpc^{-1}}$ and impose the empirical upper limit $P_{\zeta}(k_{\mathrm{Ly\alpha}})\leq 3\times 10^{-9}$ derived from Lyman-$\alpha$ forest measurements Bird et al. (2011).
Figure 5: Future prospects on the sensitivity to the amplitude of the primordial power spectrum AA. Projected constraints from galaxy stream observations are plotted in solid magenta lines corresponding to less than a 100% increase in subhalos ($\frac{N_{\rm sh}(A,k_{p},n_{b})}{N_{\rm sh}(A=0)}<2$) in four subhalo mass ranges, $10^{5}$$10^{6}$, $10^{6}$$10^{7}$, $10^{7}$$10^{8}$, and $10^{8}$$10^{9}M_{\odot}$, denoted using the lightest to darkest magenta colours respectively. Here we assume no uncertainty in the subhalo mass function for the scale-invariant power spectrum. Prospects with future $\mu$-distortion measurements with PIXIE Ünal et al. (2021) are shown with gray-dashed lines. See Fig. 4 for the Lyman-$\alpha$ bounds and expected amplitude for curvature perturbations to generate SMBH seeds.
@@
Here we denote ${\hat{k}}=k\,/\,{\rm Mpc}$ and ${\hat{k}}_{0}=1$. Current CMB $\mu$-distortion bounds are sensitive to amplitudes down to $A\sim 4\times 10^{-5}$. This probe is complementary to those DM substructures, which can be more sensitive in the mass range of $M_{\rm PBH}^{\rm seed}\gtrsim{\cal O}(10^{9})M_{\odot}$.
In each panel, the amplitude of the primordial fluctuation which produces the current SMBH energy density of $\Omega_{\textrm{SMBH}}=10^{-10}$ assuming three different statistics, $\beta_{\rm G},\beta_{\chi^{2}}$ and $\beta_{{\rm G}^{3}}$ Ünal et al. (2021) are shown with light cyan, dark cyan, and blue, respectively. Depending on the accretion efficiency $\cal A$}, the amplitude satisfying the SMBH abundance can be in between the upper (${\cal A}=1$) or lower (${\cal A}=10^{5}$) boundaries of the bands.
The sensitivity with DM structure formation is sufficient to probe non-Gaussian statistics of $\chi^{2}$ and $G^{3}$ distributions at scales $k_{p}\lesssim{\cal O}(10^{2}){\rm Mpc}^{-1}$. For $n_{b}=8$, the accessible region gets slightly narrower, however, up to $k_{p}\sim 70\,{\rm Mpc}^{-1}$ can still be probed. These regions correspond to PBH seed masses $\gtrsim 10^{9}M_{\odot}$. The Gaussian statistics case over scales of $10^{1}\lesssim k_{p}\cdot\textrm{Mpc}\lesssim 3\times 10^{4}$ for both $n_{b}$ are already excluded by measurements of the CMB $\mu$-distortion by COBE/FIRAS. For a narrower range of $30\,\textrm{Mpc}^{-1}\lesssim k_{p}\lesssim 10^{4}\textrm{\,}\textrm{Mpc}^{-1}$, COBE/FIRAS observations exclude the $\chi^{2}$ case. In all cases, the most non-Gaussian (${\rm G}^{3}$) case is not yet explored by $\mu$-distortion measurements.